Eight cells, each with an internal resistance of 0.2Ω and e.m.f of 2.2V are connected (a) in series (b) in parallel. Determine the e.m.f and the internal resistance of the batteries.
1
Expert's answer
2020-05-05T18:44:30-0400

As per the given question,

internal resistance of the cell "(r)=0.2\Omega"

Emf of each cell "(E)=2.2 V"

Number of cell "(n)=8"

a) Eight cells are connected in series, then total EMF"(E_{net})=8\times 2.2 = 17.6V"

net internal resistance of 8 cells"(r_{eq})=0.2\times 8 = 1.6\Omega"

b)

Now if all the cells are connected in parallel, then total EMF of the cell "=2.2 V"

Net resistance "=\dfrac{1}{40}=0.025\Omega"